OpenStudy (czarluc):

Exponential growth the formulas for growth y=a(1+r)^x and y=a(e^kx) where y=result a=starting amt. r=rate x=time elapsed k=constant any explanation when to use one of these???

OpenStudy (irishboy123):

these are identical if k = ln(1+r) in the second form ..... because exp(ln(1+r)) = (1+r). IOW, they should give identical results. the second form might be easier to do calculus with. the first might be easier to bung in a spreadsheet if you are using discrete time periods, eg compounding interest over months or quarters. what is really interesting that that lim r --> infinity of (1+r)^(1/r) = e. that is at the heart of why both forms work, and is the basis of the actual calculation of e itself.

OpenStudy (czarluc):

but their answers aren't identical